OpenStudy (anonymous):

What is the expression written using each base only once? 2^6 x 2^4

OpenStudy (anonymous):

4^(8+3) =4^11

OpenStudy (anonymous):

so you just add the exponets and keep the common number the same?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

pleasure

OpenStudy (asnaseer):

\[2^6*2^4=2^{6+4}=2^{10}\]